OpenStudy (anonymous):

The cost of installing insulation in a particular two-bedroom home is $1080. Present monthly heating cost are $60 but the insulation is expected to reduce heating heating costs by 10%. How many months will it take to recover the cost of the insulation?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

If you save 0.1*60 = $6 a month, then how many months will it take to reach $1080? 1080/6 = ?

OpenStudy (anonymous):

thats wht I thought is so is it 1080/6= 180?

OpenStudy (anonymous):

Yup. So it will take 180 months = 15 years to recover costs. Assuming the energy continues to cost $60 a month for the next 180 months ;)

OpenStudy (anonymous):

Thanks!